The mathematics club of RAD Senior High School organized a raffle where the prize is the latest edition of a certain brand of the calculator. the raffle aims to raise funds for their club. they were able to sell 1000 tickets that cost 5 pesos each. if they spent 3500 on the prize, what will be the expected gain of a student who will buy one ticket?

1
Expert's answer
2021-05-02T08:53:03-0400

Let "X=" gain of a student. Then


"E(X)=3500(\dfrac{1}{1000})+(-5)(\dfrac{999}{1000})=-1.495"

The expected gain of a student who will buy one ticket will be -1.495 pesos.